How do I graph?
y - 2 = -1/16(x - 1)^2

Answer by Alan3354(69443)

How do I graph?
y - 2 = -1/16(x - 1)^2
y = 2 -1/16(x - 1)^2
---
Pick values for x and find y
Plot the points
Draw a curve thru them.